Operational

  • UK-DB-1 Operational
  • US-DB-1 Operational
  • Panel Failover Operational
  • Panel Main Operational
  • LDN-ELITE-1 Operational
  • LDN-NODE-1 Operational
  • LDN-NODE-2 Operational
  • LDN-NODE-3 Operational
  • LDN-NODE-4 Operational
  • LDN-NODE-5 Operational
  • LDN-NODE-6 Operational
  • LDN-NODE-7 Operational
  • LDN-NODE-8 Operational
  • LDN-NODE-9 Operational
  • LAX-NODE-1 Operational
  • NYC-NODE-1 Operational
  • NYC-NODE-2 Operational
  • NYC-NODE-3 Operational
  • LDN-WEB-1 Operational

Incidents Incidents and scheduled downtime.

Workshop Download Problems

Resolved - 1 week ago

The Garry's Mod developers have pushed out a fix for the workshop download problems that are occurring. Please restart your servers to push the updates to your server.

Monitoring - 1 week ago

We are still monitoring the situation. Currently Steam has not pushed out a fix so we are still waiting.

Reported - 1 week ago

Steam is currently experiencing problems with their CM's after a recent maintenance. This issue is affecting most customers on London servers, however, some customers aren't being affected at all.


Currently, this is out of our hands and there isn't much we can do to resolve the problem for now until Steam push out a fix. We will continually monitor the situation and update on our status page accordingly. Thank you for your patience.

NYC Outage

Resolved - 3 weeks ago

The problem was caused by a faulty network cable. This has been replaced by on-site engineers and no incidents have been reported for around 24 hours.

Monitoring - 3 weeks ago

NYC-NODE-1 is back up and running for now. We have confirmed this was a problem on the data centre's side so we are in the process of figuring out the cause.

Outage - 3 weeks ago

NYC-NODE-1 appears to be having another outage. Upon our initial investigation we can't find a problem on our side so we believe this is another problem on the data centre's part. We will update this status post as we receive more information.

Resolved - 4 weeks ago

Datacentre has confirmed all services back up and running.

Monitoring - 4 weeks ago

The datacentre has confirmed the cause of the incident to be a power outage.

Monitoring - 4 weeks ago

Connectivity has been restored to 2/3 nodes. We will continue to monitor the situation, but NYC-NODE-1 still remains uncontactable.

Reported - 4 weeks ago

Our network provider for the New York area is experiencing network outages. We will post updates when we find out more information.